OpenStudy (anonymous):

.Which equation will result if the substitution method is used to solve the following system of linear equations? x – y = 7 x = y + 3

OpenStudy (anonymous):

Simple substitute x = y + 2 in place of y in first equation and thats ur answer

OpenStudy (studygurl14):

x - y = 7 x = y + 3 (y + 3) - y = 7

OpenStudy (anonymous):

Thanks! exactly what i thought!

OpenStudy (studygurl14):

You're welcome. :)

OpenStudy (anonymous):

what about this one? ...Solve the following system of equations by linear combination: 5x + 2y = 14 y = x – 7

OpenStudy (studygurl14):

Only one question per thread, please. But for this one, you want to eliminate a varibale in order to be able to solve the other one.

OpenStudy (anonymous):

Ok :)

OpenStudy (studygurl14):

5x + 2y = 14 y = x - 7 --> y - x = -7 -> -y + x = 7 -> x - y = 7 Now you have: 5x + 2y = 14 x - y = 7

OpenStudy (anonymous):

Thank you!

OpenStudy (studygurl14):

Looks like it will be easy to eliminate the "y" varible

OpenStudy (studygurl14):

x - y = 7 --> 2x - 2y = 14

OpenStudy (studygurl14):

Now you have: 5x + 2y = 14 2x - 2y = 14 _____________ add those two equations ^ Can you do the rest from here?

OpenStudy (anonymous):

would it be (-3,4) ?

OpenStudy (anonymous):

?? is that wrong?

OpenStudy (studygurl14):

No, sorry, but that is incorrect.